OpenStudy (anonymous):

Identities... cos^2 x = cscxcosx/tanx + cotx

OpenStudy (anonymous):

can u please recheck ur question

OpenStudy (anonymous):

cos^2x=\frac{cscx.cosx}{tanx+cotx} =\frac{\frac{1}{sinx}.cosx}{\frac{sinx}{cosx}+\frac{cosx}{sinx}} =\frac{\frac{cosx}{sinx}}{\frac{sin^2x+cos^2x}{sinx.cosx}}

OpenStudy (anonymous):